Question 25
ProbabilityThe probability that Kofi and Ama hit a target in a shooting competition are 1/6 and 1/9 respectively. What is the probability that only one of them will hit the target?
- A.1/54
- B.13/54
- C.20/27
- D.41/64
Show the answer and explanation
Answer: B (13/54)
Explanation
P(only Kofi) + P(only Ama) = (1/6)(8/9) + (5/6)(1/9) = 8/54 + 5/54 = 13/54.